OpenStudy (anonymous):

does anyone know this answer 5+4(x-8)=9-3(x-7)

OpenStudy (agent47):

5+4x-32=9-3x+21 4x+3x=30-5+32 7x=57 x=57/7

OpenStudy (anonymous):

5 + 4x - 32 = 9 - 3x + 21 7x = 30 + 32 - 5 7x = 57 And find x now..

mathslover (mathslover):

\[\huge{5+4x-32=9-3x+21}\] \[\huge{-27+4x=30-3x}\] \[\huge{-27-30=-3x-4x}\] \[\huge{-57=-7x}\] \[\huge{x=57}/7\]